OpenStudy (anonymous):

what is 6.4 x 10^-8 mi/day in feet per hour? (1 mi=5280ft) a) 8.11x10^-3ft/b b)1.408x10^-5 ft/b c)2.909x10^-10ft/b d)5.05x10^-3ft/b

OpenStudy (anonymous):

@undeadknight26

OpenStudy (anonymous):

to me it might be B but I'm not sure :/

undeadknight26 (undeadknight26):

I would think B..

OpenStudy (anonymous):

:) thx bro

OpenStudy (anonymous):

one more o_o and my last plz

undeadknight26 (undeadknight26):

Shure ladeh

OpenStudy (anonymous):

an aveergre person in the United States throws away 4.5x10^1 kg of the trash per day. In 2013, about 3.2x10^6 people in the US. About how many Kilograms of trash was thrown in the United Sation in 2013 5.3x10___ kg

undeadknight26 (undeadknight26):

I forgot how to do this stuff tbh @freckles @HelpBlahBlahBlah @Joel_the_boss @LeeEtchison Any help here mates?

OpenStudy (anonymous):

i have 5 mins left :( and its gonna be due

undeadknight26 (undeadknight26):

DO you know what these type of opperations are called?

OpenStudy (anonymous):

dude im bad at core focus : scientific notation I'm bad at it

undeadknight26 (undeadknight26):

Ohhhhh dis ish scientific notation?! I know of it but havent used it ._. And generally not in this sense....

OpenStudy (anonymous):

-.- imma put a random answer

OpenStudy (anonymous):

damn i made a 60

undeadknight26 (undeadknight26):

@arabpride @Secret-Ninja You are some of the smartest chics i know...any input?

OpenStudy (anonymous):

gtg bro its fine

undeadknight26 (undeadknight26):

Sorry mate D: Im horrible myself at scientific notations within this sense... Honestly never learned it like this XD

OpenStudy (secret-ninja):

To write a number in scientific notation: Put the decimal after the first digit and drop the zeroes. To find the exponent count the number of places from the decimal to the end of the number.

OpenStudy (secret-ninja):

Thats all I know. :/ (Because I googled it) Sorry!

OpenStudy (arabpride):

Maths is ~ ughz - idek -im too yuggh tired- good luck in stuffs xP

OpenStudy (secret-ninja):

I didn't know @arabpride was a girl. :p You always had some random picture of a baby so lol idk. XD

OpenStudy (arabpride):

LOLZ "random picture of a baby" was my nephew cx and i have NO clue why everyone thinks imma d00d ~

OpenStudy (arabpride):

@Secret-Ninja

OpenStudy (secret-ninja):

Lol XD I gtg bye peeps. :)

OpenStudy (freckles):

what is 6.4 x 10^-8 mi/day in feet per hour? (1 mi=5280ft) \[6.4 \cdot 10^{-8} \frac{miles}{day} \cdot \frac{5280 ft}{1 miles} \cdot \frac{1 day}{ 24 hours}\] since 5280ft=1mi and 1day=24 hours

OpenStudy (freckles):

the miles unit will cancel and you will have ft on top and the day unit will will cancel and you will have hours on bottom

OpenStudy (freckles):

you just got to play with the multiplication and division there

OpenStudy (freckles):

for the number part

OpenStudy (freckles):

\[\frac{6.4 \cdot 5280 }{24} \times 10^{-8} \frac{ ft}{hour}\]

OpenStudy (freckles):

oh there is another question about trash i would set up a ratio

OpenStudy (freckles):

\[\frac{x kg}{3.2 \times 10^6 people}=\frac{4.5 \times 10 kg}{1 people} \\ x (kg-people)= 4.5 \times 10 \cdot 3.2 \times 10^6 (kg-people) \\ x=4.5 \cdot 3.2 \cdot 10^{1+6}\] the fraction was setup as x amount of trash from 3.2*10^6 people is equal to 4.5*10 amount of trash from 1 person

OpenStudy (freckles):

just so you know that x was found kg/day you will need to convert that to kg/year :) since it said thrown out for the year 2013